.The probability that a bulb produced by a factory will fuse after certain period of time is 0.05. Find the probability that out of 5 such bulbs

a.None fuses

b.Not more than 2 fuse

c.More than 2 fuse
1
Expert's answer
2012-03-30T07:05:45-0400
a. None fuses

P(None fuses) = (1-0.05)^5 = 0.95^5 = 0.7738
&
b. Not more than 2 fuse

P(Not more than 2 fuse) = P(0 fuses) + P(1 fuses) + P(2 fuses) = ...
P(0 fuses) = 0.95^5
P(1 fuses) = 0.05*(1-0.05)^4
P(2 fuses) = 0.05^2*(1-0.05)^3
... = 0.95^5 + 0.05*(1-0.05)^4 + 0.05^2*(1-0.05)^3 = 0.8167.
&
c. More than 2 fuse

P(More than 2 fuse) = 1 - P(Not more than 2 fuse) = 1-0.8167 = 0.1833.